OpenStudy (anonymous):

h/h-1 - h-1/h = 1/12

OpenStudy (jhannybean):

\[\frac{h}{h-1} -\frac{h-1}{h} = -\frac{1}{12}\]ON the left side, find a common denominator.

OpenStudy (jhannybean):

What can that be? :)

OpenStudy (anonymous):

idk

OpenStudy (jhannybean):

Well, we have \(h-1\) and \(h\)

OpenStudy (anonymous):

h?

OpenStudy (jhannybean):

That can't be it, because how will we deal with the \(-1\)?

OpenStudy (jhannybean):

the easiest way to go about this is to multiply both of them together :)\[h(h-1)\]

OpenStudy (anonymous):

you mean LCM?

OpenStudy (jhannybean):

Yeah. LCM, or common denominator, :P

OpenStudy (anonymous):

h^2 - h^2 + 1 -2h?

OpenStudy (anonymous):

over h^2 - h?

OpenStudy (jhannybean):

Let's see

OpenStudy (uri):

Or..cross multiply.

OpenStudy (jhannybean):

Are you understanding what I am doing? :)

OpenStudy (anonymous):

yes i am

OpenStudy (jhannybean):

I missed something actually.

OpenStudy (jhannybean):

If we multiply al the denominators together, our LCD would be \(12(h(h-1))\)

OpenStudy (jhannybean):

Now we have a common denominator. Sorry about that.

OpenStudy (jhannybean):

Do you see why? x_x

OpenStudy (jhannybean):

If we just multiplied \(h(h-1)\) with eachother, we could be neglecting he 12 on the other side of the equation, and we cannot do that if we are to evaluate for x.

OpenStudy (jhannybean):

we would*

OpenStudy (anonymous):

what?

OpenStudy (jhannybean):

@ChiefArnav Which part have you gotten up to? Let's continue from there, step by step,

OpenStudy (jhannybean):

That way you can reference how I solved the problem and check your answer.

OpenStudy (liamschumm):

CarlosAgusta, please refrain from drawing or sending messages that are off-topic.

OpenStudy (anonymous):

\[h/(h-1) - (h-1)/h = -1/12\] So the first thing you want to do is get rid of the denominators, by multiplying everything by (h-1). This leaves us with h - ((h-1)(h-1))/h = -(h-1)/12 Now we want to get rid of the h as a denominator, by multiplying everything by h, which leaves us with: h^2 - (h-1)(h-1) = -h(h-1)/12 We can simplify this by h^2 - (h^2 -2h +1) = h(h-1)/12 All I have done there is multiply out the brackets Simplify: 2h-1 = -h(h-1)/12 Times by 12: 24h-12 = -h(h-1) Simplify: 24h-12 = -h^2 + h rearrange: 0 = -h^2 - 23h +12 Use quadratic formula with a= -1 b= -23 c= +12 I got two answers: 0.51 (approx) and -23.5 Hope this helps :D